# GNN-Transformer Cooperative Architecture for Trustworthy Graph Contrastive Learning

### v2

@@ Introduction
Compared with traditional supervised learning, self-supervised learning eliminates the dependence on labels. As one of the most representative methods of self-supervised learning, contrastive learning has been widely applied in the graph domain, i.e., Graph Contrastive Learning (GCL). GCL utilizes multiple views and positive/negative pairs for node/graph representations (Zhu et al. 2020; Peng et al. 2020; Zhu, Sun, and Koniusz 2021; Yin et al. 2022; Liu et al. 2024), the performance of which are influenced by augmentation strategies, graph encoders and loss functions.
@@
Figure 1: Augmentations on image ((a)) keep the semantic information, while augmentations on graphs ((b)) change the underlying semantic information.
@@
Most of the existing GCL methods exploit GNNs as graph encoders (Velickovic et al. 2019; Hassani and Khasahmadi 2020; Mo et al. 2022; Lee, Lee, and Park 2022; Shen et al. 2023). GNNs iteratively propagate, transform, and aggregate representations from topological neighbors to progressively update node features (Kipf and Welling 2017; Hamilton, Ying, and Leskovec 2017; Zhang et al. 2022a). With multi-layer stacking, these local features are progressively integrated to form a comprehensive perception of the entire graph structure. While this mechanism is particularly effective in handling local dependencies, especially in capturing both direct and indirect relation between nodes, it fails to capture complex global structures and long-range dependencies within graphs (Xu et al. 2019; Garg, Jegelka, and Jaakkola 2020). In contrast, Graph Transformers (GTs) are capable of effectively capturing global dependencies between nodes with self-attention mechanisms (Veličković et al. 2017; Yun et al. 2019; Kreuzer et al. 2021). GTs can handle long-range dependencies within graphs, rather than being confined to local neighborhoods. By calculating attention scores between nodes, GTs update node representations with all other nodes in the graph, thereby effectively integrating global information. This mechanism enables GTs to model complex interactions and long-term dependencies between nodes. Therefore, they are more flexible and accurate in handling dependencies between distant nodes (Yun et al. 2019), which is ignored by existing GCL methods in choosing graph encoders to learn node/graph representations. However, the quadratic complexity of GTs limits the scalability. Currently, some works introduce linear attention mechanisms to improve scalability for large-scale graphs (Wu et al. 2022, 2024; Liang, Chen, and Liang 2024).
Existing GCL methods mostly use InfoNCE as their loss functions (Zhu et al. 2020; Zhang et al. 2022b, 2023a), where each anchor point has only one positive pair. Therefore, when calculating the loss, nodes from the same class as the anchor point are pushed away. This may lead to ignorance of the potential information from similar nodes when handling node feature similarity tasks (Shen et al. 2023). While this design effectively encourages the model to distinguish between positive and negative pairs, it may lead to insufficient aggregation of nodes from the same class, thereby affecting the ability to represent graph data and capture graph structures.
@@
Graph Contrastive Learning. Current GCL methods can be categorized into three mainstream paradigms: DGI framework (Velickovic et al. 2019), InfoNCE framework (Oord, Li, and Vinyals 2018) and BGRL framework (Thakoor et al. 2022).
@@
Inspired by Deep InfoMax (Hjelm et al. 2019), DGI learns node representations by maximizing the mutual information between global and local features from the graph. MVGRL (Hassani and Khasahmadi 2020) extends DGI with graph diffusion, driving further advancement in unsupervised GCL. In spite of the competitive performance, global features in these methods may be insufficient to retain node-level embedding information.
@@
Inspired by the significant success of InfoNCE, GRACE (Zhu et al. 2020) learns node representations by pulling the same node together and pushing different nodes apart under two augmented views. This principle is widely used in both node classification (Zhu et al. 2021; Zhang et al. 2023b; Shen et al. 2023) and graph classification (You et al. 2020; Hassani and Khasahmadi 2020; Wei et al. 2023) tasks. While these methods achieve significant success, they still suffer from sampling bias issues.
To address the sampling bias problem, BGRL removes negative pairs with an exponential moving average method. AFGRL (Lee, Lee, and Park 2022) extends BGRL (Thakoor et al. 2022) with the augmentation-free strategy and node sampling technique. However, owning to the randomness of $k$-means, the sampling bias is introduced.
@@
Graph Augmentation. The existing graph augmentation strategies include node dropping (You et al. 2020), edge perturbation (Qiu et al. 2020; Zhang et al. 2023a), attribute masking (Zhu et al. 2021; Zhang et al. 2022b) and subgraph extraction (Hassani and Khasahmadi 2020). GRACE (Zhu et al. 2020) employs random edge perturbation and node feature masking to generate two views. GCA (Zhu et al. 2021) proposes an adaptive augmentation strategy which integrates both structural and attribute information. Similarly, GCS (Wei et al. 2023) utilizes the structural and semantic information to achieve adaptive graph augmentation. Furthermore, CI-GCL (Tan et al. 2024) proposes a community invariant GCL framework to maintain graph community structure. Despite the considerable success, bias is introduced due to the disturbance of the semantic.
@@
Figure 2: The architecture of GTCA. Given a graph, the GNN encoder $f_{\theta}$ and Transformer encoder $g_{\varphi}$ generate node embeddings $\bm{H}_{\theta}$ and $\bm{H}_{\varphi}$. We apply two node embeddings to obtain $k$-NNs of $v_{i}$, i.e., $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, intersects which with the topological $k$-NNs $\mathbf{T}_{i}$ to obtain positive and negative pairs for node $v_{i}$ separately, i.e., $\mathbf{P}_{i}$ and $\mathbf{N}_{i}$. Finally, we employ the contrastive loss to achieve that the anchor nodes close to positive pairs and far from negative pairs.
@@
In this section, we provide a detailed description of GTCA, including graph encoders, node sampling and contrastive loss function. Figure  2 shows the model architecture of GTCA.
@@
We employ GCN (Kipf and Welling 2016) as encoder to capture the local structural information of the nodes. GCN utilizes a series of graph convolution layers to aggregate information from the neighbors. We update each layer according to the following equation:
@@
where $\phi(\cdot)$ denotes a kernel function and $\mathbf{q}_{i}=W_{Q}^{(l)}\mathbf{h}_{i}^{(l)}$, $\mathbf{k}_{j}=W_{K}^{(l)}\mathbf{h}_{j}^{(l)}$ and $\mathbf{v}_{j}=W_{V}^{(l)}\mathbf{h}_{j}^{(l)}$. We can obtain $\bm{H}_{\varphi}$ with NodeFormer.
@@
$r\left(\hat{h}_{\mathcal{M}}\circ\hat{g}_{\mathcal{M}}\right)-r\left(\hat{h}_{\mathcal{N}}\circ\hat{g}_{\mathcal{N}}\right)$
$\leq\gamma_{\mathcal{G}}(\mathcal{M},\mathcal{N})+8L\mathfrak{R}_{N}(\mathcal{H}\circ\mathcal{G}_{\mathcal{M}})+\frac{4C}{\sqrt{N}}+2C\sqrt{\frac{2\ln(2/\delta)}{N}}$ (3)
@@
We compute the cosine similarity matrix of $\bm{H}_{\theta}$ and $\bm{H}_{\varphi}$ separately to obtain the $k$-NN sets. Specifically, we obtain the $k$-NN node set $\mathbf{B}_{i}^{\theta}$ of node $v_{i}$ with the cosine similarity between $\bm{h}_{i}^{\theta}$ and $\bm{H}_{\theta}$. Similarly, we obtain the $k$-NN node set $\mathbf{B}_{i}^{\varphi}$ for node $v_{i}$ with the cosine similarity between $\bm{h}_{i}^{\varphi}$ and $\bm{H}_{\varphi}$. Furthermore, we compute the topological $k$-NN matrix via nodes sorting according to the number of hops from other nodes to node $v_{i}$ in ascending order, resulting in the topological $k$-NN set $\mathbf{T}_{i}$ for node $v_{i}$. Finally, we obtain the positive set $\mathbf{P}_{i}$ for node $v_{i}$ from the intersection of $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$:
$\mathbf{P}_{i}=\mathbf{B}_{i}^{\theta}\cap\mathbf{B}_{i}^{\varphi}\cap\mathbf{T}_{i}$ (5)
@@
$\mathbf{N}_{i}=\mathcal{V}\setminus(\mathbf{P}_{i}\cup\{v_{i}\})$ (6)
Compared with existing augmentation based contrastive learning models, GTCA has several advantages. First, GTCA discards random or heuristic augmentation strategies which may disturb the graph structure. Second, in contrast to existing GCL methods that utilize shared-weight GNNs, GTCA uses GNN and Transformer encoders to learn diverse graph feature representations.
@@
According to (Oord, Li, and Vinyals 2018), the InfoNCE is a lower bound of the true Mutual Information (MI). Recent GCL methods use InfoNCE as the loss function (Zhu et al. 2020; Zhang et al. 2023a; Guo et al. 2023; Yu et al. 2024). Figure 3 gives an example of the InfoNCE loss. For a given anchor node, InfoNCE uses only one node to generate a positive pair and utilizes all the other nodes to generate negative pairs. This may lead to ignorance of the potential information of similar nodes.
@@
To address this issue, we design a novel contrastive loss function. Let $\bm{h}_{i}^{\theta}$ and $\bm{h}_{i}^{\varphi}$ denote the $\ell_{2}$-normalized GCN and NodeFormer embeddings of $v_{i}$ respectively. Select $\bm{h}_{i}^{\theta}$ as the anchor node, its positives come from three sources: (1) its NodeFormer embedding $\bm{h}_{i}^{\varphi}$; (2) GCN embeddings $\{\bm{h}_{j}^{\theta}\mid v_{j}\in\mathbf{P}_{i}\}$; (3) NodeFormer embeddings $\{\bm{h}_{j}^{\varphi}\mid v_{j}\in\mathbf{P}_{i}\}$. That is, the total number of positive pairs associated with node $v_{i}$ is $12|\mathbf{P}_{i}|+$ 1. Ultimately, the contrastive loss function is as follows:
@@
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, encoders $f_{\theta}$, $g_{\varphi}$ and a contrastive loss function $\mathcal{L}$ defined in Equation (11). $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$ are $k$-NNs of node $v_{i}$, $i=1,2,\cdots,N$ under two node representation views and one topology structure view. Equation (5) generates the most trustworthy positive pairs for GCL.
@@
It is obvious that the intersection set contains the minimum number of positive pairs with specific $\mathbf{B}_{i}^{\theta}$, $\mathbf{B}_{i}^{\varphi}$, and $\mathbf{T}_{i}$. The numerator of Equation (7) sums the exponential calculation with respect to $\mathbf{P}_{i}$, whereas the denominator sums the exponential calculation with respect to both $\mathbf{P}_{i}$ and $\mathbf{N}_{i}$. Therefore, when calculating $\mathbf{P}_{i}$ with Equation (5), we get the lower bound of the numerator, and also the upper bound of the contrastive loss in Equation (7). When we minimize the upper bound of the loss in Equation (11), trustworthy GCL can be achieved. ∎
@@
Algorithm 1 GTCA
Input: The adjacency matrix $\bm{A}$, the feature matrix $\bm{X}$, and the number of training epochs $J$.
Output: Feature matrix $\bm{H}$
1:  for epoch in 1 to $J$ do
2:     Generate GNN embeddings $\bm{H}_{\theta}$ and NodeFormer embeddings $\bm{H}_{\varphi}$ with GNN encoder $f_{\theta}$ , NodeFormer encoder $g_{\varphi}$, adjacency matrix $\bm{A}$ and feature matrix $\bm{X}$
[…]
4:     Calculate positive pairs $\mathbf{P}_{i}$ and negative pairs $\mathbf{N}_{i}$, $i=1,\cdots,N$ with Equation (5) and Equation (6)
5:     Compute loss $\mathcal{L}$ with Equation (11)
6:     Apply gradient descent to minimize $\mathcal{L}$ and update parameters
7:  end for
8:  Calculate the final output feature matrix with Equation (12)
9:  Return $\bm{H}$
@@
We compare GTCA with 12 state-of-the-art methods, including 2 semi-supervised GNNs, i.e., GCN (Kipf and Welling 2016) and GAT (Veličković et al. 2017), 2 semi-supervised GCL methods, i.e., CGPN (Wan et al. 2021b) and CG3 (Wan et al. 2021a), and 8 self-supervised GCL methods, i.e., DGI (Velickovic et al. 2019), GMI (Peng et al. 2020), MVGRL (Hassani and Khasahmadi 2020), GRACE (Zhu et al. 2020), GCA (Zhu et al. 2021), SUGRL (Mo et al. 2022), AFGRL (Lee, Lee, and Park 2022) and NCLA (Shen et al. 2023).
@@
We employ the above methods for node classification. For Cora dataset, we follows (Yang, Cohen, and Salakhudinov 2016) to randomly select 20 nodes per class for training, 500 nodes for validation, and the remaining nodes for testing. For Wiki-CS, Coauthor CS, Amazon-Computers and Amazon Photo datasets, we follows (Liu, Gao, and Ji 2020) to randomly select 20 nodes per class for training, 30 nodes per class for validation, and the remaining nodes for testing. We perform 20 random splits of training, validation, and testing on each dataset and report the average performance of all algorithms across these splits. All experiments are implemented in PyTorch and conducted on a server with NVIDIA GeForce 3090 (24GB memory each). Table 2 shows the hyperparameter settings of GTCA on 5 datasets.
@@
Table 4 presents the ablation study results of GTCA and its variants without $\mathbf{T}_{i}$ and with different graph encoders including GNNs $f_{\theta},f_{\varphi}$ and NodeFormers $g_{\theta},g_{\varphi}$. We observe that there is an obvious decline in performance without topology structure view on 5 datasets. This indicates that topological information is essential for enhancing node classification accuracy. In terms of graph encoders, compared with GTCA, the use of either GCN-GCN or NodeFormer-NodeFormer results in inferior performance, which validates the reasonability of the module design.
@@
Figure 4 shows that with a single graph encoder, correct ratio of positive pairs is lower than 40%. With both GCN and NodeFormer as graph encoders, the correct ratio of positive pairs varies from 30% to 60%. When GTCA exploits intersection of the $k$-NN neighborhood of GCN, Nodeformer node representation and topology structure views, the correct ratio of positive pairs exceeds 80%. This highlights the necessity and effectiveness of the current graph encoder design.
@@
Sensitivity Analysis
Figure 5 illustrates the node classification accuracy of GTCA on Cora and Amazon-Photo datasets with various hyperparameters, respectively. The optimal range of $k$ is between 300 and 600. When $k$ is too small, GTCA has inferior performance due to a scarcity of positive pairs. Conversely, a large $k$ introduces too much redundant information. Thus, it is necessary to search a proper $k$. In addition, increasing embedding dimension $E$ improve classification accuracy. Moreover, the hyperparameter $\lambda$ is crucial for model performance. Initially, an increase in $\lambda$ improves node classification accuracy. However, when $\lambda$ exceeds 0.7, the accuracy will decrease.

### v4

@@ Introduction
Compared with traditional supervised learning, self-supervised learning eliminates the dependence on labels. As one of the most representative methods of self-supervised learning, contrastive learning has been widely applied in the graph domain, i.e., Graph Contrastive Learning (GCL). GCL utilizes multiple views and positive/negative pairs for node/graph representations (Zhu et al. 2020; Peng et al. 2020; Zhu, Sun, and Koniusz 2021; Yin et al. 2022; Liu et al. 2023), the performance of which are influenced by augmentation strategies, graph encoders and loss functions.
@@
Figure 1: Augmentations on image (a) keep the semantic information, while augmentations on graphs (b) change the underlying semantic information.
@@
Most of the existing GCL methods exploit GNNs as graph encoders (Veličković et al. 2019; Hassani and Khasahmadi 2020; Mo et al. 2022; Lee, Lee, and Park 2022; Shen et al. 2023). GNNs iteratively propagate, transform, and aggregate representations from topological neighbors to progressively update node features (Kipf and Welling 2017; Hamilton, Ying, and Leskovec 2017; Zhang et al. 2022a). With multi-layer stacking, these local features are progressively integrated to form a comprehensive perception of the entire graph structure. While this mechanism is particularly effective in handling local dependencies, especially in capturing both direct and indirect relation between nodes, it fails to capture complex global structures and long-range dependencies within graphs (Xu et al. 2019; Garg, Jegelka, and Jaakkola 2020). In contrast, Graph Transformers (GTs) are capable of effectively capturing global dependencies between nodes with self-attention mechanisms (Veličković et al. 2018; Yun et al. 2019; Kreuzer et al. 2021). GTs can handle long-range dependencies within graphs, rather than being confined to local neighborhoods. By calculating attention scores between nodes, GTs update node representations with all other nodes in the graph, thereby effectively integrating global information. This mechanism enables GTs to model complex interactions and long-term dependencies between nodes. Therefore, they are more flexible and accurate in handling dependencies between distant nodes (Yun et al. 2019), which is ignored by existing GCL methods in choosing graph encoders to learn node/graph representations. However, the quadratic complexity of GTs limits the scalability. Currently, some works introduce linear attention mechanisms to improve scalability for large-scale graphs (Wu et al. 2022, 2024; Liang, Chen, and Liang 2024).
Existing GCL methods mostly use InfoNCE as their loss functions (Zhu et al. 2020; Zhang et al. 2022b, 2023), where each anchor point has only one positive pair. Therefore, when calculating the loss, nodes from the same class as the anchor point are pushed away. This may lead to ignorance of the potential information from similar nodes when handling node feature similarity tasks (Shen et al. 2023). While this design effectively encourages the model to distinguish between positive and negative pairs, it may lead to insufficient aggregation of nodes from the same class, thereby affecting the ability to represent graph data and capture graph structures.
@@
Graph Contrastive Learning. Current GCL methods can be categorized into three mainstream paradigms: DGI framework (Veličković et al. 2019), InfoNCE framework (Oord, Li, and Vinyals 2018) and BGRL framework (Thakoor et al. 2022).
@@
DGI aggregates the features of all nodes in the graph to obtain a global feature, then maximizes the mutual information between the global feature and the node features to learn node representations. Based on this framework, MVGRL (Hassani and Khasahmadi 2020) further advances the development of unsupervised graph contrastive learning through the adoption of graph diffusion and subgraph sampling methods. In spite of the competitive performance, global features in these methods may be insufficient to retain node-level embedding information.
@@
GRACE (Zhu et al. 2020) first utilizes the InfoNCE loss to maximize the mutual information between positive pairs under two augmented views, while minimizing the mutual information between negative pairs to learn node representations. This principle is widely used in both node classification (Zhu et al. 2021; Zhang et al. 2023; Shen et al. 2023) and graph classification (You et al. 2020; Hassani and Khasahmadi 2020; Wei et al. 2023) tasks. While these methods achieve significant success, they still suffer from sampling bias issues.
BGRL uses the BYOL (Grill et al. 2020) method to remove negative pairs, thereby reducing computational complexity. However, it relies on graph augmentation to obtain positive pairs, which may disturb the underlying semantics of the graph. AFGRL (Lee, Lee, and Park 2022) extends BGRL (Thakoor et al. 2022) with an augmentation-free strategy and utilizes $k$-means for positive pair sampling. However, the randomness of $k$-means may influence the performance of down-stream tasks.
@@
Graph Augmentation. The existing graph augmentation strategies include node dropping (You et al. 2020), edge perturbation (Qiu et al. 2020; Zhang et al. 2023), attribute masking (Zhu et al. 2021; Zhang et al. 2022b) and subgraph extraction (Hassani and Khasahmadi 2020). GRACE (Zhu et al. 2020) employs random edge perturbation and node feature masking to generate two views. GCA (Zhu et al. 2021) proposes an adaptive augmentation strategy which integrates both structural and attribute information. Similarly, GCS (Wei et al. 2023) utilizes the structural and semantic information to achieve adaptive graph augmentation. Furthermore, CI-GCL (Tan et al. 2024) proposes a community invariant GCL framework to maintain graph community structure. Despite the considerable success, bias is introduced due to the disturbance of the semantic.
@@
Figure 2: The architecture of GTCA. Given a graph, the GNN encoder $f_{\theta}$ and Transformer encoder $g_{\varphi}$ generate node embeddings $\bm{H}_{\theta}$ and $\bm{H}_{\varphi}$. We apply two node embeddings to obtain $k$-NNs of $v_{i}$, i.e., $\mathcal{B}_{i}^{\theta}$, $\mathcal{B}_{i}^{\varphi}$, intersects which with the topological $k$-NNs $\mathcal{T}_{i}$ to obtain positive and negative pairs for node $v_{i}$ separately, i.e., $\mathcal{P}_{i}$ and $\mathcal{N}_{i}$. Finally, we employ the contrastive loss to achieve that the anchor nodes close to positive pairs and far from negative pairs.
@@
In this section, we provide a detailed description of GTCA, including graph encoders, node sampling and contrastive loss function. Figure 2 shows the model architecture of GTCA.
@@
We employ GCN (Kipf and Welling 2017) as an encoder to capture the local structural information of the nodes. GCN utilizes a series of graph convolution layers to aggregate information from the neighbors. We update each layer according to the following equation:
@@
where $\phi(\cdot)$ denotes a kernel function and $\mathbf{q}_{i}=\bm{W}_{Q}^{(l)}\mathbf{h}_{i}^{(l)}$, $\mathbf{k}_{j}=\bm{W}_{K}^{(l)}\mathbf{h}_{j}^{(l)}$ and $\mathbf{v}_{j}=\bm{W}_{V}^{(l)}\mathbf{h}_{j}^{(l)}$. We can obtain $\bm{H}_{\varphi}$ with NodeFormer.
@@
$r\left(\hat{h}_{\mathcal{M}}\circ\hat{g}_{\mathcal{M}}\right)-r\left(\hat{h}_{\mathcal{N}}\circ\hat{g}_{\mathcal{N}}\right)\leq$
$\gamma_{\mathcal{G}}(\mathcal{M},\mathcal{N})+8L\mathfrak{R}_{N}(\mathcal{H}\circ\mathcal{G}_{\mathcal{M}})+\frac{4C}{\sqrt{N}}+2C\sqrt{\frac{2\ln(2/\delta)}{N}}$ (3)
@@
We compute the cosine similarity matrix of $\bm{H}_{\theta}$ and $\bm{H}_{\varphi}$ separately to obtain the $k$-NN sets. Specifically, we obtain the $k$-NN node set $\mathcal{B}_{i}^{\theta}$ of node $v_{i}$ with the cosine similarity between $\bm{h}_{i}^{\theta}$ and $\bm{H}_{\theta}$. Similarly, we obtain the $k$-NN node set $\mathcal{B}_{i}^{\varphi}$ for node $v_{i}$ with the cosine similarity between $\bm{h}_{i}^{\varphi}$ and $\bm{H}_{\varphi}$. Furthermore, we compute the topological $k$-NN matrix via nodes sorting according to the number of hops from other nodes to node $v_{i}$ in ascending order, resulting in the topological $k$-NN set $\mathcal{T}_{i}$ for node $v_{i}$. Finally, we obtain the positive set $\mathcal{P}_{i}$ for node $v_{i}$ from the intersection of $\mathcal{B}_{i}^{\theta}$, $\mathcal{B}_{i}^{\varphi}$, and $\mathcal{T}_{i}$:
$\mathcal{P}_{i}=\mathcal{B}_{i}^{\theta}\cap\mathcal{B}_{i}^{\varphi}\cap\mathcal{T}_{i}$ (5)
@@
$\mathcal{N}_{i}=\mathcal{V}\setminus(\mathcal{P}_{i}\cup\{v_{i}\})$ (6)
Compared with existing augmentation-based contrastive learning models, GTCA has several advantages. First, GTCA discards random or heuristic augmentation strategies which may disturb the graph structure. Second, in contrast to existing GCL methods that utilize shared-weight GNNs, GTCA uses GNN and Transformer encoders to learn diverse graph feature representations.
@@
According to (Oord, Li, and Vinyals 2018), the InfoNCE is a lower bound of the true Mutual Information (MI). Recent GCL methods use InfoNCE as the loss function (Zhu et al. 2020; Zhang et al. 2023; Guo et al. 2023; Yu et al. 2024). Figure 3 gives an example of the InfoNCE loss. For a given anchor node, InfoNCE uses only one node to generate a positive pair and utilizes all the other nodes to generate negative pairs. This may lead to ignorance of the potential information of similar nodes.
@@
To address this issue, we design a novel contrastive loss function. Let $\bm{h}_{i}^{\theta}$ and $\bm{h}_{i}^{\varphi}$ denote the $\ell_{2}$-normalized GCN and NodeFormer embeddings of $v_{i}$ respectively. Select $\bm{h}_{i}^{\theta}$ as the anchor node, its positives come from three sources: (1) its NodeFormer embedding $\bm{h}_{i}^{\varphi}$; (2) GCN embeddings $\{\bm{h}_{j}^{\theta}\mid v_{j}\in\mathcal{P}_{i}\}$; (3) NodeFormer embeddings $\{\bm{h}_{j}^{\varphi}\mid v_{j}\in\mathcal{P}_{i}\}$. That is, the total number of positive pairs associated with node $v_{i}$ is $12|\mathcal{P}_{i}|+$ 1. Ultimately, the contrastive loss function is as follows:
@@
Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, encoders $f_{\theta}$, $g_{\varphi}$ and a contrastive loss function $\mathcal{L}$ defined in Equation (11). $\mathcal{B}_{i}^{\theta}$, $\mathcal{B}_{i}^{\varphi}$, and $\mathcal{T}_{i}$ are $k$-NNs of node $v_{i}$, $i=1,2,\cdots,N$ under two node representation views and one topology structure view. Equation (5) generates the most trustworthy positive pairs for GCL.
Algorithm 1 GTCA
Input: The adjacency matrix $\bm{A}$, the feature matrix $\bm{X}$, and the number of training epochs $J$
Output: Feature matrix $\bm{H}$
1:  for epoch in 1 to $J$ do
2:     Generate GNN embeddings $\bm{H}_{\theta}$ and NodeFormer embeddings $\bm{H}_{\varphi}$ with GNN encoder $f_{\theta}$ , NodeFormer encoder $g_{\varphi}$, adjacency matrix $\bm{A}$ and feature matrix $\bm{X}$;
3:     Generate GNN $k$-NN node set $\mathcal{B}_{i}^{\theta}$, Nodeformer $k$-NN node set $\mathcal{B}_{i}^{\varphi}$ and topological $k$-NN set $\mathcal{T}_{i}$ with $\bm{H}_{\theta}$ , $\bm{H}_{\varphi}$ and adjacency matrix $\bm{A}$;
4:     Calculate positive pairs $\mathcal{P}_{i}$ and negative pairs $\mathcal{N}_{i}$, $i=1,\cdots,N$ with Equation (5) and Equation (6);
5:     Compute loss $\mathcal{L}$ with Equation (11);
6:     Apply gradient descent to minimize $\mathcal{L}$ and update parameters;
7:  end for
8:  Calculate the final output feature matrix with Equation (12);
9:  teturn $\bm{H}$ for downstream tasks;
@@
It is obvious that the intersection set contains the minimum number of positive pairs with specific $\mathcal{B}_{i}^{\theta}$, $\mathcal{B}_{i}^{\varphi}$, and $\mathcal{T}_{i}$. The numerator of Equation (7) sums the exponential calculation with respect to $\mathcal{P}_{i}$, whereas the denominator sums the exponential calculation with respect to both $\mathcal{P}_{i}$ and $\mathcal{N}_{i}$. Therefore, when calculating $\mathcal{P}_{i}$ with Equation (5), we get the lower bound of the numerator, and also the upper bound of the contrastive loss in Equation (7). When we minimize the upper bound of the loss in Equation (11), trustworthy GCL can be achieved. ∎
@@
We compare GTCA with 12 state-of-the-art methods, including 2 semi-supervised GNNs, i.e., GCN (Kipf and Welling 2017) and GAT (Veličković et al. 2018), 2 semi-supervised GCL methods, i.e., CGPN (Wan et al. 2021b) and CG3 (Wan et al. 2021a), and 8 self-supervised GCL methods, i.e., DGI (Veličković et al. 2019), GMI (Peng et al. 2020), MVGRL (Hassani and Khasahmadi 2020), GRACE (Zhu et al. 2020), GCA (Zhu et al. 2021), SUGRL (Mo et al. 2022), AFGRL (Lee, Lee, and Park 2022) and NCLA (Shen et al. 2023).
@@
We employ the above methods for node classification. For Cora dataset, we follow (Yang, Cohen, and Salakhudinov 2016) to randomly select 20 nodes per class for training, 500 nodes for validation, and the remaining nodes for testing. For Wiki-CS, Coauthor-CS, Amazon-Computers and Amazon-Photo datasets, we follow (Liu, Gao, and Ji 2020) to randomly select 20 nodes per class for training, 30 nodes per class for validation, and the remaining nodes for testing. We perform 20 random splits of training, validation, and testing on each dataset and report the average performance of all algorithms across these splits. All experiments are implemented in PyTorch and conducted on a server with NVIDIA GeForce 3090 (24GB memory each). Table 2 shows the hyperparameter settings of GTCA on 5 datasets.
@@
Table 4 presents the ablation study results of GTCA and its variants without $\mathcal{T}_{i}$ and with different graph encoders including GNNs $f_{\theta},f_{\varphi}$ and NodeFormers $g_{\theta},g_{\varphi}$. We observe that there is an obvious decline in performance without topology structure view on 5 datasets. This indicates that topological information is essential for enhancing node classification accuracy. In terms of graph encoders, compared with GTCA, the use of either GCN-GCN or NodeFormer-NodeFormer results in inferior performance, which validates the reasonability of the module design. Meanwhile, we can observe that $\text{GTCA}\_{g_{\theta},g_{\varphi}}$ ranks recond on the Coauthor-CS dataset, which has the largest number of nodes among the 5 datasets, due to the fact that Transformer-based models can effectively capture long-range dependencies.
@@
Figure 4 shows that with a single graph encoder, correct ratio of positive pairs is lower than 40%. With both GCN and NodeFormer as graph encoders, the correct ratio of positive pairs varies from 30% to 60%. When GTCA exploits intersection of the $k$-NN neighborhood of GCN, Nodeformer node representation and topology structure views, the correct ratio of positive pairs exceeds 80%. This highlights the soundness of the current graph encoder design.
@@
Sensitivity Analysis
Figure 5 illustrates the node classification accuracy of GTCA on Cora and Amazon-Photo datasets with various hyperparameters, respectively. The optimal range of $k$ is between 300 and 600. When $k$ is too small, GTCA has inferior performance due to a scarcity of positive pairs. Conversely, a large $k$ introduces too much redundant information. Thus, it is necessary to search for a proper $k$. In addition, increasing embedding dimension $E$ improves classification accuracy. Moreover, the hyperparameter $\lambda$ is crucial for model performance. Initially, an increase in $\lambda$ improves node classification accuracy. However, when $\lambda$ exceeds 0.7, the accuracy will decrease.